# A Systematic Literature Review of Parameter-Efficient Fine-Tuning for Large Code Models

### v2

@@ 5. Future Directions for PEFT Applications in Software Engineering
Building on the findings of our systematic literature review, we identify several future research directions to advance the application of PEFT techniques in SE-related activities. These directions synthesize insights across generative and non-generative tasks, model architectures, PEFT strategies, and observed evaluation gaps. We organize them into five actionable steps designed to guide both researchers and practitioners in the field.
@@
Step 1: Identify Underexplored and Emerging SE Tasks: While PEFT has seen wide application in tasks like Code Summarization and Generation, Several SE tasks remain underrepresented in the current literature. These include, but are not limited to, Software Testing, Documentation Generation, Code Quality Assessment, API Recommendation, Automated Refactoring, and Requirements Analysis. As PEFT research matures, there is ample opportunity to expand into additional task categories and workflows that are critical to real-world software development. Future work should target these areas, especially those involving complex or cross-functional requirements. Cross-task adaptation also emerges as a promising direction, where a single PEFT configuration generalizes across multiple tasks, reflecting real-world development scenarios more effectively.
@@
What to report: Clearly define the SE task and justify its importance. If proposing a novel or emerging task, document its scope, challenges, and dataset preparation methodology to aid reproducibility. Discuss why PEFT is suitable and note potential for cross-task generalization.
@@
Step 2: Align Base Models with Task Requirements: Our findings emphasize the prevalence of encoder-decoder architectures, with CodeT5 and CodeLlama variants showing dominant usage, particularly in generative tasks. Larger models like StarCoder-15B and DeepSeek-Coder-33B are gaining adoption but remain underutilized. There is also an opportunity to experiment with hybrid or task-agnostic model architectures that can support a range of SE tasks under a unified PEFT configuration.
@@
What to report: Justify model selection based on task characteristics. Include details on architecture type (encoder-only, decoder-only, encoder-decoder), model scale, and pre-training corpus. If hybrid or ensemble models are used, explain their complementary advantages.
@@
Step 3: Expand and Diversify PEFT Techniques: LoRA has emerged as the most adopted PEFT method (used in 35 studies), especially for generative tasks. Adapter-based methods dominate in non-generative settings. However, many promising PEFT approaches—such as AdaLoRA, BitFit, Prefix Tuning, and P-Tuning remain significantly under-evaluated. Future research should not only expand coverage to these methods but also explore hybrid strategies (e.g., LoRA + Prefix Tuning) and configurations tailored to task type. Furthermore, quantization-aware variants such as QLoRA demonstrate the feasibility of scaling large models efficiently and warrant deeper investigation.
@@
What to report: Detail the PEFT methods used, including training configurations, memory reduction techniques (e.g., 8-bit/4-bit quantization), and note whether any hybrid or novel PEFT combinations were tested. Clearly state why a particular method was chosen for a given task and architecture.
@@
What to report: Include evaluation metrics that match the task (e.g., BLEU, pass@1 for generative tasks; F1, MRR, AUC for non-generative tasks). Report memory and latency improvements over full fine-tuning, and include results of generalizability studies, ablations, and cross-task as well as cross-lingual evaluations.
Step 5: Promote Reproducibility and Benchmarking Platforms: Despite progress in tasks like Code Generation and Clone Detection, inconsistency in dataset usage and reporting limits comparability. Adapter-based studies, in particular, often lack detailed configuration descriptions. Future work should adopt standardized datasets and benchmarking protocols, share full implementation details, and follow community practices for reproducibility. Benchmarking diverse PEFT techniques side by side on shared tasks would help identify the most effective and efficient approaches.
@@
Together, these five steps outline a comprehensive roadmap for advancing the application of PEFT in software engineering. By broadening the scope of SE tasks, aligning base models with task demands, diversifying PEFT techniques, expanding evaluation criteria, and promoting reproducibility, future research can address current limitations while laying the groundwork for more robust, efficient, and generalizable fine-tuning solutions. We believe that systematically adopting these practices will not only accelerate progress in the field, but also ensure that PEFT methods are better positioned to meet the diverse and evolving needs of real-world software development.
@@ 6. Threats to Validity
Construct Validity relates to how accurately we identified and categorized core elements such as software engineeering tasks, model architectures, and PEFT methods. To mitigate the risk of misclassification, our approach is grounded in well-established software engineering practices relevant to the field (e.g., Watson et al. (Watson et al., 2022)) and is developed according to the guidline by Kitchenham et al. (Kitchenham et al., 2009). Additionally, multiple authors conducted independent reviews to ensure the accuracy and completeness of the task and method selection.
@@ 8. Conclusions
In this article, we conducted a systematic literature review on the application of PEFT techniques in software engineering tasks, guided by Kitchenham’s established methodology. Through a rigorous multi-phase study selection process, we identified 27 primary studies covering 17 diverse SE tasks, which we grouped into generative and non-generative categories.

### v3

@@ 6. Future Directions for PEFT Applications in Software Engineering
Building on the findings of our systematic literature review, we identify several future research directions to advance the application of PEFT techniques in SE-related activities. These directions synthesize insights across generative and non-generative tasks, model architectures, PEFT strategies, and observed evaluation gaps. We organize them into six actionable steps designed to guide both researchers and practitioners in the field.
@@
Step 1: Identify Underexplored and Emerging SE Tasks: While PEFT has seen wide application in tasks like Code Summarization and Generation, Several SE tasks remain underrepresented in the current literature. These include, but are not limited to, Software Testing, Documentation Generation, Code Quality Assessment, API Recommendation, Automated Refactoring, and Requirements Analysis. As PEFT research matures, there is ample opportunity to expand into additional task categories and workflows that are critical to real–world software development. Future work should target these areas, especially those involving complex or cross–functional requirements. Cross–task adaptation also emerges as a promising direction, where a single PEFT configuration generalizes across multiple tasks, reflecting real–world development scenarios more effectively.
@@
What to report: Clearly define the SE task and justify its importance. If proposing a novel or emerging task, document its scope, challenges, and dataset preparation methodology to aid reproducibility. Discuss why PEFT is suitable and note potential for cross–task generalization.
@@
Step 2: Align Base Models with Task Requirements: Our findings emphasize the prevalence of encoder–decoder architectures, with CodeT5 and CodeLlama variants showing dominant usage, particularly in generative tasks. Larger models like StarCoder–15B and DeepSeek–Coder–33B are gaining adoption but remain underutilized. There is also an opportunity to experiment with hybrid or task–agnostic model architectures that can support a range of SE tasks under a unified PEFT configuration.
@@
What to report: Justify model selection based on task characteristics. Include details on architecture type (encoder–only, decoder–only, encoder–decoder), model scale, and pre–training corpus. If hybrid or ensemble models are used, explain their complementary advantages.
@@
Step 3: Expand and Diversify PEFT Techniques: LoRA has emerged as the most adopted PEFT method (used in 36 studies), especially for generative tasks. Adapter-based methods dominate in non-generative settings. However, many promising PEFT approaches–such as AdaLoRA, BitFit, Prefix Tuning, and P-Tuning remain significantly under-evaluated. Additionally, advanced LoRA variants including LoRA-pro and LoRA-GA remain unexplored in software engineering contexts, despite their potential for enhanced parameter efficiency and gradient-aware optimization, respectively. Given LoRA’s demonstrated $\sim$77% success rate against full fine-tuning baselines in our analysis, these sophisticated variants could potentially address performance limitations observed in certain software engineering tasks, particularly where traditional PEFT methods showed mixed results. Future research should not only expand coverage to these methods but also explore hybrid strategies (e.g., LoRA + Prefix Tuning) and configurations tailored to task type. Furthermore, quantization-aware variants such as QLoRA demonstrate the feasibility of scaling large models efficiently and warrant deeper investigation.
@@
What to report: Detail the PEFT methods used, including training configurations, memory reduction techniques (e.g., 8–bit/4–bit quantization), and note whether any hybrid or novel PEFT combinations were tested. Clearly state why a particular method was chosen for a given task and architecture.
@@
What to report: Include evaluation metrics that match the task (e.g., BLEU, pass@1 for generative tasks; F1, MRR, AUC for non–generative tasks). Report memory and latency improvements over full fine–tuning, and include results of generalizability studies, ablations, and cross–task as well as cross–lingual evaluations.
Step 5: Promote Reproducibility and Benchmarking Platforms: Despite progress in tasks like Code Generation and Clone Detection, inconsistency in dataset usage and reporting limits comparability. Adapter–based studies, in particular, often lack detailed configuration descriptions. Future work should adopt standardized datasets and benchmarking protocols, share full implementation details, and follow community practices for reproducibility. Benchmarking diverse PEFT techniques side by side on shared tasks would help identify the most effective and efficient approaches.
@@
Step 6: Address Domain-Specific Knowledge Acquisition Limitations: Emerging evidence reveals a critical gap in current PEFT research–the fundamental challenge of adapting to domain-specific knowledge not present in pre-training data (Wang et al., 2025a). PEFT methods face substantial performance degradation when applied to private codebases or specialized domains, highlighting the limitation of updating only parameter subsets for comprehensive knowledge acquisition. Future research should prioritize developing PEFT variants specifically designed for domain adaptation scenarios, including techniques that can effectively encode new knowledge while maintaining parameter efficiency. This includes exploring progressive adaptation strategies, domain-aware parameter selection methods, and hybrid approaches that combine PEFT efficiency with targeted full parameter updates for critical knowledge domains.
What to report: Clearly distinguish between evaluations on public and private datasets, and, where applicable, report domain shift metrics–such as Jensen–Shannon divergence (Lin, 1991) or Kullback–Leibler divergence (Kullback and Leibler, 1951)–to quantify distribution differences. Complement performance results with an analysis of knowledge acquisition effectiveness, and document the specific domain characteristics and adaptation challenges considered.
Together, these six steps outline a comprehensive roadmap for advancing the application of PEFT in software engineering. By broadening the scope of SE tasks, aligning base models with task demands, diversifying PEFT techniques, expanding evaluation criteria, and promoting reproducibility, future research can address current limitations while laying the groundwork for more robust, efficient, and generalizable fine-tuning solutions. We believe that systematically adopting these practices will not only accelerate progress in the field, but also ensure that PEFT methods are better positioned to meet the diverse and evolving needs of real-world software development.
@@ 7. Threats to Validity
Construct Validity relates to how accurately we identified and categorized core elements such as software engineering tasks, model architectures, and PEFT methods. To mitigate the risk of misclassifications, our approach is grounded in well-established software engineering practices relevant to the field (e.g., Watson et al. (Watson et al., 2022)) and is developed according to the guidelines by Kitchenham et al. (Kitchenham et al., 2009). Additionally, multiple authors conducted independent reviews to ensure the accuracy and completeness of the task and method selection.
We also recognize that some practical uses of PEFT–such as those embedded in open-source repositories–may not be explicitly documented in peer-reviewed publications. While these cases may reflect real-world adoption, they often lack the methodological transparency and empirical evaluation required for systematic analysis. For instance, enabling PEFT through configuration toggles in frameworks like HuggingFace Transformers is rarely accompanied by sufficient detail on how the method was integrated, assessed, or compared against baselines. Including such undocumented or implicit usage would conflict with our review’s goal of synthesizing empirically validated and reproducible studies.
@@ 9. Conclusions
In this article, we conducted a systematic literature review on the application of PEFT techniques in software engineering tasks, guided by Kitchenham’s established methodology (Kitchenham et al., 2009). Through a rigorous multi-phase study selection process, we identified 28 primary studies covering 19 diverse SE tasks, which we grouped into generative and non-generative categories.